OpenStudy (anonymous):

find x if { -4+ (-1+2+3+.......+x)} = 437

OpenStudy (gorv):

(-1+2+3+.....+x=437+4 add 2 on both side 2-1+2+3+.....+x=441+2=443 1+2+3+....+x=443 sum on first x natunarl no= x*(x+1)/2=443 x(x+1)=443*2=886 x^2+x-886=0 solve the quadratic